OpenStudy (anonymous):

In air at 20 , sound travels at a constant speed of 344 . The light speed is equal to 299792458 . The light year is defined as the distance light travels in 1 year. Use this information to determine how many meters there are in 1 light year. Also, How many meters does light travel in 1 nanosecond?

OpenStudy (3psilon):

Use dimensional analysis We'll call the speed of light \[3x10^{8}\] so.\[\frac{ 3x10^{8}meters }{ 1 \sec }*\frac{ 60 \sec }{ 1 \min} *\frac{ 60 \min }{ 1 hr }*\frac{ 24 hrs }{ 1 day }\]

OpenStudy (3psilon):

Every thing should cancel out but meters and days. Then just multiply by 365 days to get how many meters it travels in a year

OpenStudy (anonymous):

Thanks the answer is 9.46 x 10^15 m

OpenStudy (3psilon):

np :)

OpenStudy (anonymous):

Do you know How many meters does light travel in 1 nanosecond?